# Instruction Mining: When Data Mining Meets Large Language Model Finetuning

### v2

@@ 5.2 Robustness
6 Related Work
Instruction evaluation.  The field has experienced growth with the publication of numerous instruction datasets (Taori et al., 2023; Köpf et al., 2023; Honovich et al., 2022). Chung et al. (2022) first combined multiple datasets to augment both the quantity and diversity of instruction data, achieving notable performance gains. Some newest works suggest that enhancing instruction diversity can also significantly improve instruction tuning performance (Iyer et al., 2023; Wang et al., 2023; 2022b; Longpre et al., 2023). Meanwhile, Gunasekar et al. (2023) have demonstrated that an increased proportion of high-quality data can yield enhanced performance. Other works focused on estimating the quality of prompts, Chen et al. (2023) use prompting to LLM API as an auto-grader of data quality, Gonen et al. (2022) use perplexity for prompt selection.
@@ Appendix A Search Procedure
Appendix B Indicator Descriptive Analysis
@@
Appendix C Empirical Test of Instruction Quality Evaluation Hypothesis
@@
Appendix D Other Emergent Phenomena
@@
Appendix E More Details on Parameter Fitting
@@
Appendix F Case Study
@@
F.1 Rule Selected Examples
@@
F.2 Language Model Response Examples
@@
table but he are not sure, and then they asked me to looking for it, she’s say, "Can you?",
and I responds with, "Maybe, but ain’t no sure," and he not heard me, and, "What?",
he asks, "Did you found it?".
Assistant: "Sure, I can help you with that. Here’s the corrected paragraph:
"She didn’t remember where her purse was, so I think it’s in the car but he’s not sure, and then
@@
and he didn’t hear me, and then he asked, ’What?’, he asked, ’Did you find it?’" I made the following
@@
1. Added a comma after "car" to separate the two independent clauses in the sentence.
2. Replaced "ain’t" with "are not sure" to improve the sentence’s grammar and vocabulary.
3. Added a comma after "he’s not sure" to separate the two items in the list.
4. Added a comma after "she said" to separate the quotation marks from the rest of the sentence.
5. Added a comma after "he asked" to separate the two items in the list.
6. Added a period after "he asked" to indicate the end of the sentence".
@@
User: Embrace the role of Sheldon from "The Big Bang Theory" as we delve into our conversation.
Don’t start with phrases like "As Sheldon". Let’s kick things off with the following question:
"What is your opinion on hand dryers?"
@@ 2. RNA (Ribonucleic acid) is formed from DNA and carries the genetic information to
James D. Watson in their 1953 paper, "Molecular Structure of Nucleic Acids: A Structure for
Deoxyribose Nucleic Acid." They proposed that DNA is the genetic material that carries
@@ 5. Day 5: Review the key points of the Opium Wars and discuss how the drama, mime, or theater
Appendix G Random Sampling Experiment Results
@@
Table 21: Random experiment results 2.

### v3

# Instruction Mining: Instruction Data Selection for Tuning Large Language Models

@@ Appendix A Search Procedure
Appendix B Related Work
Instruction data management.  The field has experienced growth with the publication of numerous instruction datasets (Taori et al., 2023; Köpf et al., 2023; Honovich et al., 2022). Chung et al. (2022) first combined multiple datasets to augment both the quantity and diversity of instruction data, achieving notable performance gains. Some newest works suggest that enhancing instruction diversity can also significantly improve instruction tuning performance (Iyer et al., 2023; Wang et al., 2023; 2022b; Longpre et al., 2023). Wu et al. (2023) purpose DiverseEvol, a methodology that iteratively enhances the diversity of the training subset to optimize performance. Other works focused on the quality of prompts, Gunasekar et al. (2023) have demonstrated that an increased proportion of high-quality data can yield enhanced performance. Chen et al. (2023) use prompting to LLM API as an auto-grader of data quality, Gonen et al. (2022) use perplexity for prompt selection. Considering both quality and diversity, Bukharin & Zhao (2023) purposed Quality-Diversity Instruction Tuning to control dataset diversity and quality. Meanwhile, Engstrom et al. (2024) select data by modeling the learning algorithm’s utilization of training data for predicting target tasks.
Appendix C Justification of Loss as a Suitable Indicator
In this section, we provide further evidence for selecting loss as a suitable quality indicator in our method. To effectively select high-quality data examples without sacrificing the total spent time, the selected indicator should (1) be able to represent data quality to some extent (2) not take too much time or expense during quality inference. The loss indicator we used in this work is computed on a golden evaluation set which is of relatively high quality. One risk for using loss as the indicator is that by minimizing the distance between the selected dataset and golden evaluation set, the finetuned model can overfit on the golden set, which means that the finetuned model can only perform well on instructions which are similar to the ones in the golden set. However, our experiment results presented in Section 4 reveal that the finetuned model did not overfit on the selected golden set. The effectiveness appears to be robust over different models and datasets.
To further investigate whether loss can be representative of data quality or model performance, we conduct experiments to see how inference loss value varies with different models. Results are presented in Table 7.
Model Name Model Size Loss(self-instruct) Loss(mt-bench) mt-bench score
LLaMA-2 7B 1.270 1.025 1.28
Vicuna-v1.5 7B 1.075 0.869 6.17
LLaMA-2-chat 7B 1.031 0.798 6.27
Table 7: Inference loss values for different models.
In addition, we also discover that the computed loss value, mt-bench scores and OpenLLM benchmark scores can sometimes present different trends. This phenomenon is similar to the one proposed by Zeng et al. (2023). Our hypothesis is that metrics like OpenLLM benchmark scores and MMLU are more focused on question answering, while metrics like mt-bench and loss are more focused on the model’s ability to interact with humans. This difference might be the cause of the indicators’ presenting different judgements sometimes.
Appendix D Indicator Descriptive Analysis
@@
Appendix E Empirical Test of Instruction Quality Evaluation Hypothesis
@@
Appendix F Other Emergent Phenomena
@@
Appendix G More Details on Parameter Fitting
@@
Appendix H Case Study
@@
H.1 Rule Selected Examples
@@
H.2 Language Model Response Examples
@@
table but he are not sure, and then they asked me to looking for it, she’s say, ”Can you?”,
and I responds with, ”Maybe, but ain’t no sure,” and he not heard me, and, ”What?”,
he asks, ”Did you found it?”.
Assistant: ”Sure, I can help you with that. Here’s the corrected paragraph:
”She didn’t remember where her purse was, so I think it’s in the car but he’s not sure, and then
@@
and he didn’t hear me, and then he asked, ’What?’, he asked, ’Did you find it?’” I made the following
@@
1. Added a comma after ”car” to separate the two independent clauses in the sentence.
2. Replaced ”ain’t” with ”are not sure” to improve the sentence’s grammar and vocabulary.
3. Added a comma after ”he’s not sure” to separate the two items in the list.
4. Added a comma after ”she said” to separate the quotation marks from the rest of the sentence.
5. Added a comma after ”he asked” to separate the two items in the list.
6. Added a period after ”he asked” to indicate the end of the sentence”.
@@
User: Embrace the role of Sheldon from ”The Big Bang Theory” as we delve into our conversation.
Don’t start with phrases like ”As Sheldon”. Let’s kick things off with the following question:
”What is your opinion on hand dryers?”
@@ 2. RNA (Ribonucleic acid) is formed from DNA and carries the genetic information to
James D. Watson in their 1953 paper, ”Molecular Structure of Nucleic Acids: A Structure for
Deoxyribose Nucleic Acid.” They proposed that DNA is the genetic material that carries
@@ 5. Day 5: Review the key points of the Opium Wars and discuss how the drama, mime, or theater
Appendix I Random Sampling Experiment Results
@@
Table 23: Random experiment results 3.